# TalkingHeadBench: A Multi-Modal Benchmark & Analysis of Talking-Head DeepFake Detection

### v1

@@ B.1.2 Data Curation and Quality Control
B.1.3 Train/Test Splits and Real Data Integration
@@ B.2 Computational Resources
B.3 Hyperparameters and Training Details for Detectors
@@
Appendix C Detailed Per-Generator Artifact Observations
@@ C.3 Hallo2 [8]
  • Static Peripheral Hair Artifacts: Hair situated outside the primary head bounding box frequently remained static during head motion, creating a visually jarring "detached" effect—a specific type of hair artifact and temporal inconsistency.
@@
C.4 AniPortrait [46]
C.4.1 Audio-driven
@@ C.4.2 Video-driven
  • Catastrophic "Head Explosion" Artifacts: A frequent failure mode involved severe face warping/distortion, where the synthesized head appeared to disintegrate or become overlaid with disparate visual elements (often misidentified hands or other body parts) from the driving video, a critical spatial inconsistency.
@@ C.5 EMOPortraits [12]
  • "Floating Head" Artifacts: Facial and head modifications were often restricted to the cephalic region, leading to a dissociation from neck and body movements. Significant corporeal motion in the driving video could result in the head appearing detached or "floating," a clear example of inconsistent head movement/pose and spatial inconsistency.
@@
Appendix D Additional Experimental Results
D.1 MAGI-1 Commercial Generator Results
We present detection performance on deepfakes generated by the MAGI-1 commercial model. This model was used as an additional unseen generator to test detector robustness. Since MAGI-1 does not have audio output, we only use it to evaluate the performance of CADDM, TALL, and DeepFake-Adapter (DF-Adapter). At T1 (Tab. D.9), both DeepFake-Adapter and TALL exhibit excellent performance on MAGI-1. When the threshold is tightened to T0.1 (Tab. D.10), DF-Adapter’s performance drops (T1=0.89) while TALL retains an excellent score (T1=1.00). This observation, particularly TALL’s strong performance and DF-Adapter’s drop at stricter thresholds, is consistent with trends observed with other generators, offering insights into detector generalizability against unseen commercial models.
Table D.9: Protocol 1 — MAGI-1 (Test Only). Best numbers for Magi-1 are bold. T1 metric performance colored (Excellent: >0.99, Good: 0.95-0.99, Reasonable: 0.85-0.95, Fair: 0.75-0.85, Poor: <0.75).
Detector AUC T1 T0.1
CADDM 0.91 0.55 0.55
TALL 1.00 1.00 1.00
DF-Adapter 1.00 1.00 1.00
Table D.10: Protocol 3 — MAGI-1 (Test Only). Best numbers per (generator) are bold
Detector AUC T1 T0.1
CADDM 0.94 0.52 0.52
TALL 1.00 1.00 1.00
DF-Adapter 1.00 1.00 0.89
Appendix E Additional Explainability Results
@@
E.1 Grad-CAM Visualizations for TALL [50]
Figure E.7: Grad-CAM visualizations of the TALL detector on correctly classified, high-confidence fake samples across Protocols 1–3 (one sample per generator). TALL predominantly focuses on facial features such as the eyes, nose, and mouth, but strategically shifts its attention to background or peripheral regions for certain generators. In Protocol 1, where the model has seen all generators during training, it focuses on background cues for Hallo2 and EMOPortraits, suggesting that these regions carry distinctive generative artifacts. In Protocol 3, the detector localizes to the neck region for EMOPortraits, while maintaining facial focus elsewhere. These examples reflect TALL’s adaptive attention and its ability to leverage generator-specific cues that contribute to its superior generalization and detection performance.
Figure E.7 presents Grad-CAM visualizations of the TALL detector across Protocols 1–3, using one correctly classified high-confidence fake sample per generator per protocol. These examples reflect cases where TALL confidently and accurately identifies manipulations, allowing us to interpret what visual evidence it relies on for detection.
@@ E.1 Grad-CAM Visualizations for TALL [50]
E.2 Grad-CAM Visualizations for DeepFake-Adapter [40]
Figure E.8: Grad-CAM visualizations of DeepFake-Adapter across Protocols 1–3 (same samples as used for TALL). Compared to TALL, DF-Adapter shows broader and less localized attention maps, often covering both facial and background regions with reduced spatial focus. While it identifies meaningful areas in some cases (e.g., the mouth region in EMOPortraits under Protocol 1), attention weakens in Protocols 2 and 3, especially for EMOPortraits, correlating with its performance drop. The model appears to rely on a mixture of weak signals across the image rather than generator-specific artifact patterns, reflecting its less consistent generalization under distribution shifts.
Figure E.8 presents Grad-CAM results for the DeepFake-Adapter model across Protocols 1–3 using the same sample images previously analyzed for TALL. Compared to TALL, whose attention maps are generally well-localized around facial or generator-specific regions, DF-Adapter displays broader and more diffuse attention, often spanning both the face and background without a clearly focused region of interest.

### v2

@@ C.1 Dataset Generation, Splits, and Curation Details
Train/Test Splits and Real Data Integration
@@ C.2 Computational Resources
C.3 Hyperparameters and Training Details for Detectors
@@
Appendix D Detailed Per-Generator Artifact Observations
@@ D.3 Hallo2 [cui2024hallo2]
  • Static Peripheral Hair Artifacts: Hair situated outside the primary head bounding box frequently remained static during head motion, creating a visually jarring ”detached” effect—a specific type of hair artifact and temporal inconsistency.
@@
D.4 AniPortrait [wei2024aniportrait]
Audio-driven
@@ D.4 AniPortrait [wei2024aniportrait]
  • Catastrophic ”Head Explosion” Artifacts: A frequent failure mode involved severe face warping/distortion, where the synthesized head appeared to disintegrate or become overlaid with disparate visual elements (often misidentified hands or other body parts) from the driving video, a critical spatial inconsistency.
@@ D.5 EMOPortraits [drobyshev2024emoportraits]
  • ”Floating Head” Artifacts: Facial and head modifications were often restricted to the cephalic region, leading to a dissociation from neck and body movements. Significant corporeal motion in the driving video could result in the head appearing detached or ”floating,” a clear example of inconsistent head movement/pose and spatial inconsistency.
@@
Appendix E Additional Experimental Results
E.1 MAGI-1 Commercial Generator Results
We present detection performance on deepfakes generated by the MAGI-1 commercial model, an unseen generator included to test detector robustness against proprietary systems. Since MAGI-1 does not provide audio output, we restrict evaluation to detectors operating on visual input: CADDM, HiFi-Net, TALL, and DeepFake-Adapter (DF-Adapter). Results are reported in Tab. 7.
Overall, TALL achieves perfect scores across all metrics (AUC, T1, T0.1), indicating strong robustness to MAGI-1’s generation pipeline. DF-Adapter also performs near-perfectly, though its performance drops slightly at stricter thresholds (e.g., T0.1 falling to 0.99 on some splits). By contrast, CADDM achieves only moderate performance (AUC 0.88–0.97, T1 around 0.3–0.6), reflecting limited generalization under MAGI-1. HiFi-Net performs worst overall, with low T1 values (<< 0.39) and near-zero T0.1 on most splits, suggesting that its learned features fail to transfer to this commercial generator.
These results highlight two key trends. First, detectors with strong multi-generator training performance (TALL, DF-Adapter) transfer well to unseen commercial generators, achieving excellent or near-excellent robustness. Second, detectors such as CADDM and HiFi-Net, which are more fragile under distribution shifts, fail to generalize effectively despite maintaining higher AUC values. This reinforces the broader insight from our benchmark: aggregate metrics such as AUC can obscure vulnerabilities, while threshold-based measures (T1, T0.1) reveal substantial gaps in robustness to emerging generators.
Table 7: Supplementary results on MAGI-1 commercial generator.
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 CADDM [dong2023implicit] 0.88 0.29 0.29 0.95 0.39 0.39 0.91 0.53 0.53 0.96 0.62 0.62 0.97 0.62 0.62 0.97 0.64 0.64 HiFi-Net [hifi_net_xiaoguo] 0.66 0.02 0.00 0.89 0.39 0.23 0.85 0.04 0.02 0.73 0.02 0.00 0.86 0.06 0.04 0.87 0.09 0.02 TALL [xu2023tall] 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 DF-Adapter [shao2025deepfake] 1.00 1.00 1.00 1.00 0.99 0.99 1.00 0.99 0.99 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00
E.2 Hallo3 Academic Generator Results
We evaluate detector robustness on Hallo3, a recent academic talking-head generator. Results are shown in Tab. 8. As with MAGI-1, we restrict evaluation to detectors operating on visual input.
TALL achieves perfect scores across all metrics (AUC, T1, T0.1), demonstrating excellent generalization to Hallo3. DF-Adapter also performs nearly perfectly, with T1 and T0.1 values consistently at 0.99–1.00 across all splits. By contrast, CADDM shows only moderate performance, with AUC ranging from 0.67 to 0.97 and T1 values spanning 0.24–0.91 depending on the split. HiFi-Net performs inconsistently: while it reaches very high scores on EMOPortraits (AUC = 1.00, T1 = 0.99), it collapses on other splits (e.g., T1 = 0.19 on AniAudio, T0.1 = 0.00 on LivePortrait), highlighting its fragility under generator and protocol shifts.
These findings mirror broader trends from our benchmark: detectors such as TALL and DF-Adapter maintain strong robustness on unseen generators, while CADDM and HiFi-Net struggle to generalize reliably. The variability in CADDM and the failure cases of HiFi-Net underscore the importance of evaluating with stricter thresholds, as aggregate metrics like AUC alone would obscure these weaknesses.
Table 8: Supplementary results on Hallo3 academic generator.
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 AUC T1 T0.1 CADDM [dong2023implicit] 0.67 0.24 0.24 0.70 0.40 0.40 0.93 0.76 0.76 0.82 0.64 0.64 0.97 0.91 0.91 0.93 0.69 0.69 HiFi-Net [hifi_net_xiaoguo] 0.84 0.19 0.18 0.88 0.30 0.27 0.91 0.24 0.20 0.65 0.24 0.12 1.00 0.99 0.99 0.87 0.02 0.00 TALL [xu2023tall] 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 DF-Adapter [shao2025deepfake] 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 0.99 0.99
Appendix F Additional Explainability Results
@@
F.1 Grad-CAM Visualizations for TALL [xu2023tall]
Figure 7: Grad-CAM visualizations of the TALL detector on correctly classified, high-confidence fake samples across Protocols 1–3 (one sample per generator). TALL predominantly focuses on facial features such as the eyes, nose, and mouth, but strategically shifts its attention to background or peripheral regions for certain generators. In Protocol 1, where the model has seen all generators during training, it focuses on background cues for Hallo2 and EMOPortraits, suggesting that these regions carry distinctive generative artifacts. In Protocol 3, the detector localizes to the neck region for EMOPortraits, while maintaining facial focus elsewhere. These examples reflect TALL’s adaptive attention and its ability to leverage generator-specific cues that contribute to its superior generalization and detection performance.
Figure 7 presents Grad-CAM visualizations of the TALL detector across Protocols 1–3, using one correctly classified high-confidence fake sample per generator per protocol. These examples reflect cases where TALL confidently and accurately identifies manipulations, allowing us to interpret what visual evidence it relies on for detection.
@@ F.1 Grad-CAM Visualizations for TALL [xu2023tall]
F.2 Grad-CAM Visualizations for DeepFake-Adapter [shao2025deepfake]
Figure 8: Grad-CAM visualizations of DeepFake-Adapter across Protocols 1–3 (same samples as used for TALL). Compared to TALL, DF-Adapter shows broader and less localized attention maps, often covering both facial and background regions with reduced spatial focus. While it identifies meaningful areas in some cases (e.g., the mouth region in EMOPortraits under Protocol 1), attention weakens in Protocols 2 and 3, especially for EMOPortraits, correlating with its performance drop. The model appears to rely on a mixture of weak signals across the image rather than generator-specific artifact patterns, reflecting its less consistent generalization under distribution shifts.
Figure 8 presents Grad-CAM results for the DeepFake-Adapter model across Protocols 1–3 using the same sample images previously analyzed for TALL. Compared to TALL, whose attention maps are generally well-localized around facial or generator-specific regions, DF-Adapter displays broader and more diffuse attention, often spanning both the face and background without a clearly focused region of interest.
@@ F.2 Grad-CAM Visualizations for DeepFake-Adapter [shao2025deepfake]
Appendix G Demographic Bias Audit
To assess the fairness of the benchmarked detectors and understand biases inherited from the source datasets (FFHQ, CelebV-HQ), we conducted a demographic audit on our test set using classifications from the FairFace model [karkkainen2021fairface]. We analyzed detector performance across three axes: race, gender, and age. Tab 9 summarized below indicate minimal bias along gender lines but reveal notable performance disparities across racial and age subgroups. This analysis highlights an important area for future work in both dataset creation and algorithmic fairness for deepfake detection.
Table 9: Detector performance by demographics. Disparities appear across race and age, while gender differences are minimal.
(a) Race
Group Accuracy
Asian 88.89%
African American 81.58%
Indian 87.23%
White 87.55%
(b) Gender
Group Accuracy
Female 87.77%
Male 86.98%
(c) Age Group
Age Group Accuracy
0–2 95.00%
3–9 93.88%
10–19 90.91%
20–29 87.90%
30–39 84.48%
40–49 82.56%
50–59 85.37%
60–69 94.12%
70+ 80.00%